OpenStudy (ashleynguyenx3):

Find the value(s) of c guaranteed by the Mean Value Theorem for Integrals for the function over the given interval. f(x)=x-2(sqrt x), interval [0,2]

OpenStudy (anonymous):

f ' (c) = [ f ( b) - f ( a) ] / ( b -a)

OpenStudy (anonymous):

@ashleynguyenx3 Where are you?

OpenStudy (ashleynguyenx3):

Sorry

OpenStudy (ashleynguyenx3):

Yes, I know.

OpenStudy (ashleynguyenx3):

I got the f(b) - F(a) part

OpenStudy (ashleynguyenx3):

Why would we need to find f'(x) I thought we'd find the anitderivative

OpenStudy (anonymous):

Are you sure you comprehend the MVT?

OpenStudy (ashleynguyenx3):

Nope, my teacher didn't really go over it.

OpenStudy (ashleynguyenx3):

My book gives me the fomula 

OpenStudy (anonymous):

Find f ' (c) Find f ( b) - f ( a) ] / ( b -a) Equating them

OpenStudy (turingtest):

@ashleynguyenx3 your formula is the same as @Chlorophyll 's if you take the derivative of both sides and use the fundamental theorem of calculus.

OpenStudy (ashleynguyenx3):

f'(c)= -2x f(b) - f(a) / (b-a) = 0

OpenStudy (ashleynguyenx3):

Wait idk I'm doing.

OpenStudy (ashleynguyenx3):

I was looking at the wrong equation. So sorry.

OpenStudy (ashleynguyenx3):

f(0) = 0 f(2) = 2- 2\[\sqrt x\]

OpenStudy (ashleynguyenx3):

f(2) = 2 - 2(sqrt 2)

OpenStudy (ashleynguyenx3):

And then f'(c) = x - 3x^(3/2)

OpenStudy (ashleynguyenx3):

1 - 3x^(3/2)

OpenStudy (ashleynguyenx3):

4x^(3/2) / 3

OpenStudy (ashleynguyenx3):

Oh I get it, I did the antiderivative.

OpenStudy (ashleynguyenx3):

So my f'(c) = 1 - 2(1/ sqrt x)

OpenStudy (ashleynguyenx3):

2 sqrt x *

OpenStudy (ashleynguyenx3):

f ' ( c) = 1 - 1/sqrt x

OpenStudy (anonymous):

Find f ' (c) Find f ( b) - f ( a) ] / ( b -a) Equating them

OpenStudy (ashleynguyenx3):

Okay, now I get it. You used the original MVT theorem.

OpenStudy (anonymous):

Find c?

OpenStudy (ashleynguyenx3):

So I just set them equal to find the value of x, well c? But I don't understand how come my book says there are two answers.

OpenStudy (turingtest):

chlorophyll and I apparently both misunderstood the question

OpenStudy (anonymous):

I see, I skip the phrase " for integral" in your post!

OpenStudy (anonymous):

So sorry about it!

OpenStudy (ashleynguyenx3):

No, I'm sorry.

OpenStudy (ashleynguyenx3):

So could you please explain it now?

OpenStudy (anonymous):

What point you don't understand in the solution?

OpenStudy (turingtest):

I don't think i understand where the factor if two goes between step 3 and 4 frankly :P

OpenStudy (ashleynguyenx3):

I don't understand after you solve the antiderivative and set it equal to f(c)(2-0)

OpenStudy (ashleynguyenx3):

I know it's the formula, but how do you solve it?

OpenStudy (anonymous):

Since the concept is slope at specific point f'(c) = [ f(b) - f(a) ]/ ( b -a) => derivative for integral at x =c is f(c) ( b -a) = f(b) - f(a)

OpenStudy (anonymous):

To be clear F' (c) = [ F ( b) - F (a) ] / ( b -a) f ( c) ( b -a ) = F ( b) - F (a)

OpenStudy (ashleynguyenx3):

Okay, I think I some what get it. From the the work that they show I just don't get how they got from f(c)(2-0)=(6 - 8 sqrt 2)/3 to c - 2 sqrt c = ( 3 -4 sqrt 2)/3

OpenStudy (anonymous):

f(x) = 2 - 2√x ->f (c) = 2 - 2√c

OpenStudy (anonymous):

Then simplify both sides for 2: -> ( 3 -4 √2)

OpenStudy (ashleynguyenx3):

Nope, thank you! :)

OpenStudy (anonymous):

Hundred % clear?

OpenStudy (ashleynguyenx3):

Not quite, but I think I'll get it eventually

OpenStudy (ashleynguyenx3):

AB, we're on a semester system so I'll be taking BC in January.

OpenStudy (anonymous):

Wow, it's tough :s

OpenStudy (ashleynguyenx3):

Not to be mean, but I don't think my teacher is a good teacher.